OpenStudy (anonymous):

(3 pt) Which property does the statement illustrate? Column A Column B 1. (p + 3) + q = p + (3 + q) 2. 8 + (9 + 4) = (8 + 9) + 4 3. 4(x – 2) = 4x – 8 4. 7(3 + 5) = 7 • 3 + 7 • 5 A. associative property of addition B. distributive property

OpenStudy (anonymous):

ill give medal

OpenStudy (mathstudent55):

Associative (change of grouping): (a + b) + c = a + (b + c) Distributive: a(b + c) = ab + ac

OpenStudy (anonymous):

which one goes with it?

OpenStudy (mathstudent55):

Look at the name of each property and the equation near it. Then compare that with the equations in your problem.

OpenStudy (anonymous):

there all the same number though

OpenStudy (anonymous):

thanks i figured it out

OpenStudy (anonymous):

thx fur helpin mah lil sis out

OpenStudy (anonymous):

@mathstudent55 i gave chu medal fur helpin her out

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

i did tooo